OpenStudy (anonymous):

Solve this system of equations by substitution. 2x+y=6 3x+2y=10 Answer: (2,2)

OpenStudy (anonymous):

y=6-2x 3x+2(6-2x)=10 3x+12-4x=10 -x+12=10 -x=10-12 -x=-2 x=2

OpenStudy (anonymous):

how do i get the other coordinate of 2?

OpenStudy (anonymous):

since x=2 2x+y=6 2(2)+y=6 4+y=6 y=6-4 y=2

OpenStudy (anonymous):

THANK YOU! :)